Find the complements of 

[(A'+B)C']' =?

[(AB'+C)D'+E']' =?

[A+((BC')'+D)]" =?


1
Expert's answer
2021-12-13T11:12:49-0500


(i) Using de Morgan's theorem


"\begin{bmatrix}
 (A'+B) &C'\\
 
\end{bmatrix}=(A'+B)'+C''"


Applying involution law

"=(A'+B)'+C"


Applying de Morgan's theorem

"=A''.B'+C"


Applying involution law

"=A.B'+C"


(ii)

Applying de Morgan's theorem

"=\begin{bmatrix}
 (AB'+C)D' 
 
\end{bmatrix}'.E''"


Applying involution law

"=\begin{bmatrix}
 a
(AB'+C) & D'
 
\end{bmatrix}'.E"


Applying de Morgan's law

"=\begin{bmatrix}
 (AB'+C)'+D'' 
 
\end{bmatrix}.E"


Applying double negation

"\begin{bmatrix}
 (AB'+C)'+D 
 
\end{bmatrix}.E"


Applying de Morgan's theorem

"=\begin{bmatrix}
 (AB')'.C'+D
 
\end{bmatrix}.E"


"=\begin{bmatrix}
 (A'+B'').C'+D
 
\end{bmatrix}.E"


"=\begin{bmatrix}
 (A'+B).C'+D 
 
\end{bmatrix}.E"


"=\begin{bmatrix}
 A'C'+BC'+D 
 
\end{bmatrix}.E"


"=E(D+C').(B+D+A')"



Applying involution law

"=\begin{bmatrix}
 A+((BC')'+D) \\
 
\end{bmatrix}''"


Using involution law

"=A+((BC')'+D)"


Applying de Morgan's theorem

"=A+(B'+C'')+D"


Applying involution law

"=A+B'+C+D"